OpenStudy (anonymous):

A bag has 5 red marbles, 6 blue marbles and 4 black marbles. What is the probability of picking a blue marble, replacing it, and then picking a black marble?

OpenStudy (kropot72):

The probability of first picking a blue marble is given by: \[P(blue)=\frac{number\ of\ blue\ marbles}{total\ number\ of\ marbles}=you\ can\ calculate\] Similarly the probability of a black marble is: \[P(black)=\frac{4}{15}\] The required probability is given by: \[P(blue) \times P(black)=you\ can\ calculate\]